Direct Jet (DJ) has a bond issue with face value of $1,000, coupon rate of 7 percent, with 10 years to maturity. The interest is paid semi-annually and the bond is currently selling at 10 percent discount. Assuming your MARR is 8% compounded semi-annually, is this a good investment?
1
Expert's answer
2020-05-04T11:59:58-0400

We should calculate NPV to evaluate if this investment is good.

"NPV = -0.9\times1,000 + \frac{70}{(1 + 0.08/2)^1} + \frac{70}{(1 + 0.08/2)^2} + ... + \frac{70 + 1000}{(1 + 0.08/2)^{20}} = -273.47."

So, as NPV < 0, then this investment is not good.